OpenStudy (gabylovesyou):

how do i find the solutions for x2 + 3x − 18 = 0 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

x^2 + 3x - 18 = 0 factors to (x+6)(x-3) = 0

jimthompson5910 (jim_thompson5910):

now use the zero product property to solve

OpenStudy (gabylovesyou):

what is that ?

jimthompson5910 (jim_thompson5910):

zero product property if A*B = 0 then A = 0 or B = 0

jimthompson5910 (jim_thompson5910):

example (x+7)(x-5) = 0 turns into x+7 = 0 or x-5 = 0

OpenStudy (gabylovesyou):

ohh ok

OpenStudy (gabylovesyou):

x + 6 = 0 or x - 3 = 0

jimthompson5910 (jim_thompson5910):

good, keep going

jimthompson5910 (jim_thompson5910):

if x+6 = 0, then x must be?

OpenStudy (gabylovesyou):

-6

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

do the same for x-3=0 to get the other solution

OpenStudy (gabylovesyou):

3

jimthompson5910 (jim_thompson5910):

perfect

jimthompson5910 (jim_thompson5910):

your two solutions are x = -6 or x = 3

OpenStudy (gabylovesyou):

thanks :)

jimthompson5910 (jim_thompson5910):

the hardest part is probably factoring huh?

jimthompson5910 (jim_thompson5910):

if so, then I recommend you practice that some more